A furnace delivers 8.0 x 10^7 cal per hour. How many kilojoules per sec is this? (1 cal = 4.184 joules). show by hand conversion
Answer #1

time = 1 hour = 3600 s

Energy = 8.0*10^7 cal

= 8.0*10^7*4.184 J

= 3.35*10^8 J

= 3.35*10^5 KJ

KJ per second = Energy / time

= (3.35*10^5 KJ)/(3600 s)

= 93 KJ/s

Answer: 93 KJ/s